# On anomalous dissipation induced by transport noise

### v2

@@ 4.1. Scaling limit with cut-off at a fixed Reynolds number
Then there exists a constant $N\geq 0$ independent of $\zeta_{0}$ such that
(4.12) $\sup_{t\in[0,\infty)}\|\zeta(t,\cdot)\|_{H^{\delta}}+\|\zeta\|_{L^{2}({\mathbb{R}}_{+};H^{1+\delta})}\leq N\|\zeta_{0}\|_{H^{\delta}}.$
@@
Let $p_{0}(\nu,\mu)>2$ be as in Proposition 4.2. Fix $p\in(2,p_{0})$ such that $1-2/p<\delta$ and $r,r_{0}$ such that $r<r_{0}<1-2/p$. Note that the choice of $p$ yields $\zeta_{0}\in B^{1-2/p}_{2,p}$, cf. (3.8). Next, by Proposition 4.5, there exists $K_{0}=K_{0}(N,r_{0})>0$ such that the unique global solution $u_{\mathrm{det}}$ to (4.10) satisfies
@@
Next, to apply Proposition 4.2 with $R=K_{0}+1$ and $(r,r_{0})$ as above, we check that (4.2) has a unique solution on $[0,1]$ and it is given by $\zeta_{{\rm det}}^{\nu}$. To begin, as $\zeta_{{\rm det}}^{\nu}$ is a unique global solution to (4.10) satisfying the bound (4.13), then it is also a local solution to (4.2) on $[0,1]$ with $\zeta_{{\rm det}}^{\nu}\in C([0,1];H^{r})\cap L^{2}(0,1;H^{1})$. It remains to discuss the uniqueness. To this end, we mimic a stopping-time argument. Let $\zeta_{{\rm det}}^{\nu,\prime}$ be another global solution to (4.2) with $\zeta_{{\rm det}}^{\nu,\prime}\in C([0,1];H^{r})$. Let
$e^{\prime}\stackrel{{\scriptstyle{\rm def}}}{{=}}\inf\{t\in[0,1]\,:\,\|\zeta_{{\rm det}}^{\nu,\prime}(t,\cdot)\|_{H^{r}}\geq R\}\quad\text{ where }\quad\inf\varnothing\stackrel{{\scriptstyle{\rm def}}}{{=}}1.$
If $e^{\prime}=1$, then $\zeta_{{\rm det}}^{\nu}=\zeta_{{\rm det}}^{\nu,\prime}$ by uniqueness of solutions to (4.10). If $e^{\prime}<1$, then still using the uniqueness of solutions to (4.10) we get $\zeta_{{\rm det}}^{\nu}=\zeta_{{\rm det}}^{\nu,\prime}$ on $[0,e^{\prime}]$. As $\sup_{t\in[0,e^{\prime}]}\|\zeta_{{\rm det}}^{\nu}(t)\|_{H^{r}}\leq\sup_{t\in[0,1]}\|\zeta_{{\rm det}}^{\nu}(t)\|_{H^{r_{0}}}\leq K_{0}$ and $R>K_{0}$ by construction, it follows that $\|\zeta_{{\rm det}}^{\nu,\prime}(e^{\prime},\cdot)\|_{H^{r}}<R$ which contradicts the definition of $e^{\prime}$ and $\zeta_{{\rm det}}^{\nu,\prime}\in C([0,1];H^{r})$. This proves $e^{\prime}=1$ and hence uniqueness of (4.2) follows from the one of solutions to (4.10).
@@
$\mathbf{P}\Big{(}\sup_{t\in[0,1]}\|\zeta_{{\rm cut}}^{\nu}-\zeta_{{\rm det}}^{\nu}\|_{H^{r}}\leq 1\Big{)}>1-\varepsilon$
where $\zeta_{{\rm cut}}^{\nu}$ is the unique global $p$-solution to (4.1) with the above choice of $(R,r)$. Set ${\Omega_{0}\stackrel{{\scriptstyle{\rm def}}}{{=}}\big{\{}\sup_{t\in[0,1]}\|\zeta_{{\rm cut}}^{\nu}-\zeta_{{\rm det}}^{\nu}\|_{H^{r}}\leq 1\big{\}}.}$ Now, to conclude, it remains to prove that
@@
and $\inf\varnothing\stackrel{{\scriptstyle{\rm def}}}{{=}}1$. Note that, by (4.13), we have
@@
Moreover, $\phi_{R,r}(\zeta_{{\rm cut}}^{\nu})=1$ on $[0,\tau_{0}]$. Hence, $(\zeta_{{\rm cut}}^{\nu},\tau_{0})$ is a local solution to the 2D NSEs (1.1). By uniqueness and maximality of the solution $\zeta_{{\rm cut}}^{\nu}$ to (1.1), we obtain
(4.16) $\zeta^{\nu}(t)=\zeta_{{\rm cut}}^{\nu}(t)\ \text{ a.s.\ for all $t\in[0,\tau_{0}]$,}$
cf., Definition 2.2. Hence, the claim (4.14) follows from (4.15) and (4.16). ∎
@@
Due to Corollary 4.4, the proof of Theorem 1.1 now follows almost verbatim from the one of Theorem 1.2 given at the end of Subsection 3.2. ∎
5. The case of 2D NSEs – Velocity formulation
@@
(5.1) $\partial_{t}u^{\nu}+(u^{\nu}\cdot\nabla)u^{\nu}&=-\nabla p^{\nu}+\nu\Delta u^{\nu}\\ &+\sqrt{2\mu}\sum_{k\in{\mathbb{Z}}^{2}_{0}}\big{[}-\nabla\widetilde{p}_{n}^{\,\nu}+\theta_{k}^{\nu}\,(\sigma_{k}\cdot\nabla)u^{\nu}\big{]}\circ\dot{W}_{t}^{k}&\text{on }&\mathbb{T}^{2},\\ \nabla\cdot u^{\nu}&=0&\text{on }&\mathbb{T}^{2},\\ u^{\nu}(0,\cdot)&=u_{0}&\text{on }&\mathbb{T}^{2}.\end{aligned}\right.$
where $(W^{k})_{k\in{\mathbb{Z}}^{2}_{0}}$ is are complex-valued Brownian motions as in Subsection 2.2. In Remark 5.2 below we also discuss the three-dimensional case. Let us point out that taking the curl $(u_{x},u_{y})\mapsto\partial_{x}u_{y}-\partial_{y}u_{x}$ in (5.1) does not give the vorticity formulation (1.1). Therefore, the results in this section are not related to Theorem 1.1 besides the proof strategy. The stochastic NSEs (5.1) can be physically derived via two-scale by mimicking the two-scale decomposition as in Subsection 1.2. The reader is referred to [28, 65, 66] for more rigorous physical derivations of (5.1).
@@
Formally, $\mathbb{P}f$ can be written as $f-\nabla\Delta^{-1}(\nabla\cdot f)$. We set =$\mathbb{Q}\stackrel{{\scriptstyle{\rm def}}}{{=}}\mathrm{Id}-\mathbb{P}.$ From standard Fourier analysis, it follows that $\mathbb{Q}$ and $\mathbb{P}$ restrict to bounded linear operators on $H^{s,q}$ and $B^{s}_{q,p}$ for $s\in{\mathbb{R}}$ and $q,p\in(1,\infty)$. Finally, we can introduce function spaces of divergence-free vector fields: For $\mathcal{A}\in\{L^{q},H^{s,q},B^{s}_{q,p}\}$, we set
@@
endowed with the natural norm. $\ell^{2}$-values function spaces are defined analogously.
@@
with initial data $u^{\nu}(0,\cdot)=u_{0}$, both on $\mathbb{T}^{2}$. Following the arguments in Subsection 2.2, we reformulate the Stratonovich noise in (5.2) as an Itô term plus a correction term. As in [37, Section 2], by (2.4), it holds that
@@
Further comments on the transformation (5.3) can be found in [9, Section 1]. Arguing as in (2.5), we can reformulate the noise terms in (5.3) and (5.4) using the real-valued coefficients $\Re\sigma_{k}$ and $\Im\sigma_{k}$. Finally, solutions to (5.1) in the form (5.2)-(5.4) can be defined as in Definition 2.2 by replacing the usual Lebesgue and Sobolev spaces by the Lebesgue and Sobolev spaces of divergence-free vector fields, i.e., $\mathbb{H}^{1}(\mathbb{T}^{2})$ and $\mathbb{L}^{2}(\mathbb{T}^{2})$, respectively. We omit the details for brevity.
@@
The claims of Theorem 1.1 holds with $\zeta_{0}$ and $\zeta^{\nu}$ replaced by a mean-free vector field $u_{0}\in\mathbb{H}^{\delta}(\mathbb{T}^{2})$ and the solution $u^{\nu}$ to the 2D NSEs (5.1), respectively; provided in (2) the condition $\mu>-\ln(1-\eta)$ is replaced by $\mu>-4\ln(1-\eta)$.
Before going into the proof of Theorem 5.1, let us discuss its physical interpretation. In the context of KB theory of 2D turbulence [13, 56, 43], anomalous dissipation of energy does not hold. Thus, the result of Theorem 5.1 is not physically relevant. Let us emphasize that this does not conflict with the physical interests of Theorem 1.1 as the latter concerns with the stochastic vorticity formulation which cannot be obtained from the velocity one, see the comments below (5.1). Arguing as in Subsection 1.2, we conjecture that this is due to the incorrect scaling for $\mathrm{supp}\,\theta^{\nu}$ in Theorem 5.1. Physically speaking, following the two-scale terminology of Subsection 1.2, the former corresponds to an overestimation of the ‘excitement’ of the small scales. Thus, we conjecture that $\lim_{\nu\downarrow 0}\nu^{1/2-\varepsilon}N^{\nu}=\infty$ for all $\varepsilon>0$ where $N^{\nu}$ satisfies $\mathrm{supp}\,\theta^{\nu}\subseteq\{k\in{\mathbb{Z}}^{2}_{0}\,:\,|k|\eqsim N^{\nu}\}$ and $(\theta^{\nu})_{\nu\in(0,1)}$ is such that anomalous dissipation of energy for (5.1) holds.
@@
A version of Theorem 1.1 also holds for 3D NSEs in velocity formulation provided, at a fixed viscosity $\nu>0$, well-posedness and a-priori bounds for the 3D NSEs hold in the sub-critical space $\mathbb{L}^{r}(\mathbb{T}^{3})$ with $r>3$. We refrain from formulating the latter assumption, as it implies the regularization by noise for 3D NSEs, which is one of the major open problems in stochastic fluid dynamics. The reader is referred to [2] for a recent result.
The proof of Theorem 5.1 is analogous to the one of Theorem 1.1, where instead of the Meyers’ estimate of Theorem A.1 one uses the one for the turbulent Stokes system A.2. Therefore, we only give a sketch here. Before doing so, let us give the following analogue of Proposition 4.5 for the 2D NSEs in the velocity formulation.
@@
The above is well-known to experts and can be obtained by using an interpolation result as in [45] and the fact that $H^{\delta}_{{\rm mz}}=(L_{{\rm mz}}^{2},H_{\rm mz}^{1})_{\delta,2}$ where $(\cdot,\cdot)_{\delta,2}$ is the real interpolation functor (see e.g., [16] for details) and the subscript ${\rm mz}$ stands for mean-zero. The case $\delta>1$ of Proposition 5.3 also holds and it follows from the case $\delta<1$ and the sub-criticality of $H^{\delta}$ with $\delta>0$ for (5.5).
Proof of Theorem 5.1 – Sketch.
The proof follows almost verbatim the one of Theorem 1.1. Note that the key estimate (4.5) also holds for the nonlinearity 2D NSEs in velocity formulation, see the comments below (4.5). The factor ‘$4$’ in the condition $\mu>-4\ln(1-\eta)$ for having a prescribed rate as in Theorem 1.1(2) comes from the (non-vanishing) behaviour as $n\to\infty$ of the Itô-correction $Q_{\theta}$ in (5.4) with $\theta=\theta^{n}$ as in (3.4), see [61, Theorem 3.1]. ∎

### v3

@@ 4.1. Scaling limit with cut-off at a fixed Reynolds number
(4.17) $\sup_{t\in[0,\infty)}\|\zeta(t,\cdot)\|_{H^{\delta}}+\|\zeta\|_{L^{2}({\mathbb{R}}_{+};H^{1+\delta})}\leq N(\|\zeta_{0}\|_{L^{2}})\,\|\zeta_{0}\|_{H^{\delta}}.$
@@
Let $p_{0}(\nu,\mu)>2$ be as in Proposition 4.2. Fix $p\in(2,p_{0})$ such that $1-2/p<\delta$ and $r,r_{0}$ such that $r<r_{0}<1-2/p$. Note that the choice of $p$ yields $\zeta_{0}\in B^{1-2/p}_{2,p}$, cf. (3.8). Next, by Proposition 4.5, there exists $K_{0}=K_{0}(N,r_{0})>0$ such that the unique global solution $\zeta_{{\rm det}}^{\nu}$ to (4.15) satisfies
@@
Next, to apply Proposition 4.2 with $R=K_{0}+1$ and $(r,r_{0})$ as above, we check that (4.2) has a unique solution on $[0,1]$ and it is given by $\zeta_{{\rm det}}^{\nu}$. To begin, as $\zeta_{{\rm det}}^{\nu}$ is a unique global solution to (4.15) satisfying the bound (4.18), then it is also a local solution to (4.2) on $[0,1]$ with $\zeta_{{\rm det}}^{\nu}\in C([0,1];H^{r})\cap L^{2}(0,1;H^{1})$. It remains to discuss the uniqueness (see the comments below Proposition 4.2 for the uniqueness concept used here). To this end, we mimic a stopping-time argument. Let $\zeta_{{\rm cut,d}}^{\nu,\prime}$ be another global solution to (4.2) with $\zeta_{{\rm cut,d}}^{\nu,\prime}\in C([0,1];H^{r})$. Let
$e^{\prime}\stackrel{{\scriptstyle{\rm def}}}{{=}}\inf\{t\in[0,1]\,:\,\|\zeta_{{\rm cut,d}}^{\nu,\prime}(t,\cdot)\|_{H^{r}}\geq R\}\quad\text{ where }\quad\inf\varnothing\stackrel{{\scriptstyle{\rm def}}}{{=}}1.$
If $e^{\prime}=1$, then $\zeta_{{\rm det}}^{\nu}=\zeta_{{\rm cut,d}}^{\nu,\prime}$ by uniqueness of solutions to (4.15). If $e^{\prime}<1$, then still using the uniqueness of solutions to (4.15) we get $\zeta_{{\rm det}}^{\nu}=\zeta_{{\rm det}}^{\nu,\prime}$ on $0,e^{\prime}]$. As $\sup_{t\in[0,e^{\prime}]}\|\zeta_{{\rm det}}^{\nu}(t)\|_{H^{r}}\leq\sup_{t\in[0,1]}\|\zeta_{{\rm det}}^{\nu}(t)\|_{H^{r_{0}}}\leq K_{0}$ and $R>K_{0}$ by construction, it follows that $\|\zeta_{{\rm cut,d}}^{\nu,\prime}(e^{\prime},\cdot)\|_{H^{r}}<R$ which contradicts the definition of $e^{\prime}$ and $\zeta_{{\rm cut,d}}^{\nu,\prime}\in C([0,1];H^{r})$. This proves $e^{\prime}=1$ and hence uniqueness of (4.2) follows from the one of solutions to (4.15).
@@
$\mathbf{P}\Big(\sup_{t\in[0,1]}\|\zeta_{{\rm cut}}^{\nu}-\zeta_{{\rm det}}^{\nu}\|_{H^{r}}\leq\varepsilon\Big)>1-\varepsilon$
where $\zeta_{{\rm cut}}^{\nu}$ is the unique global $p$-solution to (4.1) with the above choice of $(R,r)$. Set ${\Omega_{0}\stackrel{{\scriptstyle{\rm def}}}{{=}}\big\{\sup_{t\in[0,1]}\|\zeta_{{\rm cut}}^{\nu}-\zeta_{{\rm det}}^{\nu}\|_{H^{r}}\leq\varepsilon\big\}.}$ Now, to conclude, it remains to prove that
@@
and $\inf\varnothing\stackrel{{\scriptstyle{\rm def}}}{{=}}1$. Note that, by (4.18) and $\varepsilon<1$, we have
@@
Moreover, $\phi_{R,r}(\zeta_{{\rm cut}}^{\nu})=1$ on $0,\tau_{0}]$. Hence, $(\zeta_{{\rm cut}}^{\nu},\tau_{0})$ is a local solution to the 2D NSEs (1.1). By uniqueness and maximality of the global solution $\zeta^{\nu}$ to (1.1) (see the comments below Definition 2.2), we obtain
(4.21) $\zeta^{\nu}(t)=\zeta_{{\rm cut}}^{\nu}(t)\ \text{ a.s.\ for all $t\in[0,\tau_{0}]$.}$
Hence, the claim (4.19) follows from (4.20) and (4.21). ∎
@@
Due to Corollary 4.4, the proof of Theorem 1.1 now follows almost verbatim from that of Theorem 1.2 given at the end of Subsection 3.2. ∎
Similar to Proposition 3.6, from Corollary 4.4 one can deduce the following variant of Theorem 1.1.
Proposition 4.6 (Anomalous dissipation of enstrophy by transport noise – 2D NSEs II).
Let $N\geq 1$ and $\delta>0$. For all $\mu>0$ and $\eta\in(0,1-e^{-\mu/(4\pi^{2})})$, there exists a family $(\theta^{\nu})_{\nu\in(0,1)}\subseteq\mathcal{S}_{\ell^{2}}^{0}$ such that, for all mean-zero $\zeta_{0}\in H^{\delta}(\mathbb{T}^{2})$ satisfying $N^{-1}\leq\|\zeta_{0}\|_{L^{2}}$ and $\|\zeta_{0}\|_{H^{\delta}}\leq N$, it holds that
$\inf_{\nu\in(0,1)}\mathbf{E}\int_{0}^{1}\int_{\mathbb{T}^{d}}\nu\,|\nabla\zeta^{\nu}|^{2}\,\mathrm{d}x\,\mathrm{d}t\geq\frac{\eta}{2}\,\|\zeta_{0}\|_{L^{2}(\mathbb{T}^{d})}^{2},$
where $\zeta^{\nu}$ is the unique global smooth solution to (1.1).
5. Anomalous dissipation for 2D NSEs in velocity formulation
@@
(5.1) $\left\{\begin{aligned} \partial_{t}u^{\nu}+(u^{\nu}\cdot\nabla)u^{\nu}&=-\nabla p^{\nu}+\nu\Delta u^{\nu}\\ &+\sqrt{2\mu}\sum_{k\in{\mathbb{Z}}^{2}_{0}}\big[-\nabla\widetilde{p}_{n}^{\,\nu}+\theta_{k}^{\nu}\,(\sigma_{k}\cdot\nabla)u^{\nu}\big]\circ\dot{W}_{t}^{k}&\text{on }&\mathbb{T}^{2},\\ \nabla\cdot u^{\nu}&=0&\text{on }&\mathbb{T}^{2},\\ u^{\nu}(0,\cdot)&=u_{0}&\text{on }&\mathbb{T}^{2}.\end{aligned}\right$.
where $(W^{k})_{k\in{\mathbb{Z}}^{2}_{0}}$ is are complex-valued Brownian motions as in Subsection 2.2. In Remark 5.2 below, we also discuss the three-dimensional case. Let us point out that taking the curl operator $(u_{x},u_{y})\mapsto\partial_{x}u_{y}-\partial_{y}u_{x}$ in (5.1) does not give the vorticity formulation (1.1) (see [67, Remark 2.1] for details). Therefore, the results in this section are not related to Theorem 1.1 besides the proof strategy. The stochastic NSEs (5.1) can be derived by mimicking the two-scale decomposition as in Subsection 1.2. The reader is referred to [31, 72, 73] for more rigorous physical derivations of (5.1).
@@
Formally, $\mathbb{P}f$ can be written as $f-\nabla\Delta^{-1}(\nabla\cdot f)$. We set $\mathbb{Q}\stackrel{{\scriptstyle{\rm def}}}{{=}}\mathrm{Id}-\mathbb{P}.$ From standard Fourier analysis, it follows that $\mathbb{Q}$ and $\mathbb{P}$ restrict to bounded linear operators on $H^{s,q}$ and $B^{s}_{q,p}$ for $s\in{\mathbb{R}}$ and $q,p\in(1,\infty)$. Finally, we can introduce function spaces of divergence-free vector fields: For $\mathcal{A}\in\{L^{q},H^{s,q},B^{s}_{q,p}\}$, we set
@@
endowed with the natural norm. $\ell^{2}$-valued function spaces are defined analogously.
@@
with initial data $u^{\nu}(0,\cdot)=u_{0}$, both on $\mathbb{T}^{2}$. Following the arguments in Subsection 2.2, we reformulate the Stratonovich noise in (5.2) as an Itô term plus a correction term. As in [42, Section 2], by (2.4), it holds that
@@
Further comments on the transformation (5.3) can be found in [10, Section 1]. Arguing as in (2.6), we can reformulate the noise terms in (5.3) and (5.4) using the real-valued coefficients $\Re\sigma_{k}$ and $\Im\sigma_{k}$. Finally, solutions to (5.1) in the form (5.2)-(5.4) can be defined as in Definition 2.2 by interpreting the convective term $(u^{\nu}\cdot\nabla)u^{\nu}$ in the conservative form $\nabla\cdot(u^{\nu}\otimes u^{\nu})$, and replacing the usual Lebesgue and Sobolev spaces by the Lebesgue and Sobolev spaces of divergence-free vector fields, i.e., $\mathbb{H}^{1}(\mathbb{T}^{2})$ and $\mathbb{L}^{2}(\mathbb{T}^{2})$, respectively. We omit the details for brevity. For completeness, we mention that the existence of a unique global solution to (5.1) with $u_{0}\in\mathbb{L}^{2}(\mathbb{T}^{2})$ with finite energy (i.e., $u\in L^{2}(0,T;\mathbb{H}^{1}(\mathbb{T}^{2}))\cap C([0,T];\mathbb{L}^{2}(\mathbb{T}^{2}))$ for all $T<\infty$) is well-known and can be, for instance, obtained via either [8, Theorem 3.4] or [11, Theorem 7.10].
@@
Let $N\geq 1$ and $\delta>0$ be fixed. Then the following assertions hold.
  • For all $\mu>0$, there exists a family $(\theta^{\nu})_{\nu\in(0,1)}\subseteq\mathcal{S}_{\ell^{2}}^{0}$ such that, for all mean-zero $u_{0}\in\mathbb{H}^{\delta}(\mathbb{T}^{2})$ satisfying $\|u_{0}\|_{H^{\delta}}\leq N$, we have
    $\liminf_{\nu\downarrow 0}\mathbf{E}\int_{0}^{1}\int_{\mathbb{T}^{2}}\nu\,|\nabla u^{\nu}|^{2}\,\mathrm{d}x\,\mathrm{d}t\geq\frac{1}{2}(1-e^{-\mu/(16\pi^{2})})\,\|\zeta_{0}\|_{L^{2}(\mathbb{T}^{2})}^{2}.$
  • For all $\mu>0$ and $\eta\in(0,1-e^{-\mu/(16\pi^{2})})$, there exists a family $(\theta^{\nu})_{\nu\in(0,1)}\subseteq\mathcal{S}_{\ell^{2}}^{0}$ such that, for all mean-zero $u_{0}\in\mathbb{H}^{\delta}(\mathbb{T}^{2})$ satisfying $N^{-1}\leq\|u_{0}\|_{L^{2}}$ and $\|u_{0}\|_{H^{\delta}}\leq N$, it holds that
    $\inf_{\nu\in(0,1)}\mathbf{E}\int_{0}^{1}\int_{\mathbb{T}^{d}}\nu\,|\nabla u^{\nu}|^{2}\,\mathrm{d}x\,\mathrm{d}t\geq\frac{\eta}{2}\,\|u_{0}\|_{L^{2}(\mathbb{T}^{d})}^{2}.$
In the above statements, $u^{\nu}$ is the unique global smooth solution to (5.1).
The appearence of the factor $\mu/(16\pi^{2})$ in Theorem 5.1, rather than $\mu/(4\pi^{2})$ as in Theorem 1.1, is due to the asymptotic behaviour of the Itô-Stratonovich corrector $Q_{\theta}$ under the scaling limit which will be carried out in Proposition 5.3 below.
Before going into the proof of Theorem 5.1, let us discuss its physical interpretation. In the context of the KB theory of 2D turbulence [15, 62, 48], anomalous dissipation of energy does not hold. Thus, the result of Theorem 5.1 is not physically relevant. Let us emphasize that this does not conflict with the physical interests of Theorem 1.1 as the latter is concerned with the stochastic vorticity formulation, which cannot be obtained from the velocity one, see the comments below (5.1). Arguing as in Subsection 1.2, we conjecture that this is due to the incorrect scaling for $\mathrm{supp}\,\theta^{\nu}$ in Theorem 5.1. Physically speaking, following the two-scale terminology of Subsection 1.2, the former corresponds to an overestimation of the ‘excitement’ of the small scales. Thus, we conjecture that $\lim_{\nu\downarrow 0}\nu^{1/2-\varepsilon}N^{\nu}=\infty$ for all $\varepsilon>0$ where $N^{\nu}$ satisfies $\mathrm{supp}\,\theta^{\nu}\subseteq\{k\in{\mathbb{Z}}^{2}_{0}\,:\,|k|\eqsim N^{\nu}\}$ and $(\theta^{\nu})_{\nu\in(0,1)}$ is such that anomalous dissipation of energy for (5.1) holds.
@@
A version of Theorem 1.1 also holds for 3D NSEs in velocity formulation, provided, at a fixed viscosity $\nu>0$, well-posedness and a-priori bounds for the 3D NSEs hold in the sub-critical space $\mathbb{L}^{r}(\mathbb{T}^{3})$ with $r>3$. We refrain from formulating the latter assumption, as it implies the regularization by noise for 3D NSEs, which is one of the major open problems in stochastic fluid dynamics. The reader is referred to [2] for a recent result.
As for the 2D NSE in vorticity formulation (1.1), the proof of Theorem 5.1 is analogous to that of Theorem 1.1 and Proposition 3.6. As for the vorticity formulation (1.1) analyzed in Section 4, the key ingredients are the scaling limit as in Proposition 4.2 and the $H^{\delta}$-estimate of Proposition 4.5.
To extend the Proposition 4.2 to the stochastic 2D NSEs in velocity form (5.1), we consider the following truncated SPDE:
(5.5) $\left\{\begin{aligned} \partial_{t}u_{{\rm cut}}^{\nu}&+\phi_{R,r}(u_{{\rm cut}}^{\nu})\,[(u_{{\rm cut}}^{\nu}\cdot\nabla)u_{{\rm cut}}^{\nu}]=\nu\Delta u_{{\rm cut}}^{\nu}\\ &\qquad\qquad\qquad+\sqrt{2\mu}\sum_{k\in{\mathbb{Z}}^{2}_{0}}\theta_{k}\,(\sigma_{k}\cdot\nabla)u_{{\rm cut}}^{\nu}\circ\dot{W}_{t}^{k}&\text{ on }&\mathbb{T}^{2},\\ u_{{\rm cut}}^{\nu}(0,\cdot)&=u_{0}&\text{ on }&\mathbb{T}^{2},\end{aligned}\right$.
where $\theta=(\theta_{k})_{k\in{\mathbb{Z}}^{2}_{0}}\in\ell^{2}$, $\phi_{R,r}(u_{{\rm cut}}^{\nu})\stackrel{{\scriptstyle{\rm def}}}{{=}}\phi\big(R^{-1}\|u_{{\rm cut}}^{\nu}\|_{H^{r}}\big)$ and $\phi\in C^{\infty}([0,\infty))$ is a cutoff function such that $\mathrm{supp}\,\phi\subseteq[0,2]$ and $\phi=1$ on $[0,1]$.
Next, we prove an analogue of Proposition 4.2 for the 2D NSEs with transport noise in the velocity formulation (5.1).
Proposition 5.3 (Scaling limit – 2D NSEs in velocity form with cut-off).
Fix $\nu\in(0,1)$ and $\mu>0$. Let $(\theta^{n})_{n\geq 1}\subseteq\mathcal{S}_{\ell^{2}}^{0}$ be as in (3.4). Then there exists $p_{0}(\nu,\mu)>2$ for which the following assertion holds. If for $p\in(2,p_{0}]$, $r\in(0,1-2/p)$ and $R>0$ we have
  1. (1)
    $u_{0}\in\mathbb{B}^{1-2/p}_{2,p}$;
  2. (2)
    there exists a unique solution $u^{\nu}\in C([0,1];\mathbb{H}^{r})\cap L^{2}(0,1;\mathbb{H}^{1})$ on $[0,1]$ to
    (5.6) $\left\{\begin{aligned} \partial_{t}u_{\mathrm{cut,d}}^{\nu}+\phi_{R,r}(u_{\mathrm{cut,d}}^{\nu})\,[\nabla p_{\mathrm{cut,d}}^{\nu}+(u_{\mathrm{cut,d}}^{\nu}\cdot\nabla)u_{\mathrm{cut,d}}^{\nu}]&=\Big(\nu+\frac{\mu}{4}\Big)\Delta u_{\mathrm{cut,d}}^{\nu}&\text{on }&\mathbb{T}^{2},\\ u_{\mathrm{cut,d}}^{\nu}(0,\cdot)&=u_{0}&\text{on }&\mathbb{T}^{2};\end{aligned}\right$.
then for all $r_{0}\in(0,1-2/p)$ and all sequences $(u_{0,n})\subseteq\mathbb{B}^{1-2/p}_{2,p}$ such that $u_{0,n}\to u_{0}$ in $\mathbb{B}^{1-2/p}_{2,p}$ it holds that
$\lim_{n\to\infty}\mathbf{P}\Big(\sup_{t\in[0,1]}\|u_{{\rm cut},n}^{\nu}(t)-u_{\mathrm{cut,d}}^{\nu}(t)\|_{H^{r_{0}}}\geq\varepsilon\Big)=0\ \text{ for all }\varepsilon>0,$
where $u_{{\rm cut},n}^{\nu}$ is the unique global $p$-solution to (5.5) with $\theta=\theta^{n}$ (whose definition is analogous to the one in Definition 4.1) and initial data $u_{0,n}$.
There are two minor differences between Proposition 5.3 and Proposition 4.2. First, the sequence chosen for the scaling limit, $(\theta^{n})_{n\geq 1}\in\mathcal{S}_{\ell^{2}}^{0}$, is more restricted. Second, the limiting deterministic PDE (5.6) features a reduced enhanced dissipation coefficient, namely $\mu/4$. Both differences stem from the asymptotic behavior of $Q_{\theta^{n}}$ as $n\to\infty$, cf. (5.7) below.
Proof.
The proof of Proposition 5.3 closely follows that of Proposition 4.2. We therefore comment only on the key modifications. Firstly, the nonlinear term $\mathbb{P}[\nabla\cdot(u^{\nu}\otimes u^{\nu})]$ admits estimates analogous to those used in the 2D NSEs with transport noise in vorticity formulation (1.1). For instance, parallel to (4.5), for all $u\in\mathbb{H}^{1}(\mathbb{T}^{2})$, for all $r>0$, it holds that
$\|\mathbb{P}[\nabla\cdot(u\otimes u)]\|_{H^{-1}}\lesssim\|u\|_{L^{4}}^{2}\lesssim\|u\|_{H^{r}}^{1+\kappa_{r}}\|u\|_{H^{1}}^{1-\kappa_{r}}$
where $\kappa_{r}=\frac{r}{1-r}>0$. Similarly, the estimate (4.11) also holds for the nonlinearity in the velocity form. Hence, from the previous facts, one can prove an analogue of Lemma 4.3, where in corresponding Step 2, one employs the Meyers estimates for the turbulent Stokes system of Theorem A.3 instead of Theorem A.1.
Second, the factor $\frac{\mu}{4}$ follows by repeating Step 2 of Proposition 3.1 and using that, for all divergence-free vector field $\phi\in C^{\infty}(\mathbb{T}^{d};{\mathbb{R}}^{d})$ and $(\theta_{n})_{n\geq 1}$ as in (3.4) it holds that
(5.7) $\lim_{n\to\infty}Q_{\theta^{n}}(\varphi)=\frac{3}{4}\mu\,\Delta\varphi\ \text{ in }\ H^{-\delta}(\mathbb{T}^{d};{\mathbb{R}}^{d})\ \text{ for all }\delta>0;$
see [67, Theorem 3.1], and also [42, Theorem 5.1] for the three-dimensional case.
The remainder of the proof follows by repeating the arguments in Proposition 4.2 with minimal modification. ∎
Finally, similar to Proposition 4.5, the following result holds.
@@
The above is well-known to experts and can be obtained by using an interpolation result, as in [50], and the fact that $H^{\delta}_{{\rm mz}}=(L_{{\rm mz}}^{2},H_{\rm mz}^{1})_{\delta,2}$ where $(\cdot,\cdot)_{\delta,2}$ is the real interpolation functor (see e.g., [18] for details) and the subscript $\rm mz$} stands for mean-zero. The case $\delta>1$ of Proposition 5.4 also holds and it follows from the case $\delta<1$ and the sub-criticality of $H^{\delta}$ with $\delta>0$ for (5.8).
Proof of Theorem 5.1.
By Propositions 5.3 and 5.4, the conclusion of Corollary 4.4 remains valid when $\zeta^{\nu},\zeta_{{\rm cut,d}}^{\nu}$ and $\zeta_{{\rm det}}^{\nu}$ are replaced by $u^{\nu},u_{\mathrm{cut,d}}^{\nu}$ and $u_{\mathrm{det}}^{\nu}$, respectively; where $u_{\mathrm{det}}^{\nu}$ is the unique global solution to (5.8) with $\kappa=\nu+\mu/4$. Consequently, the assertions of Theorem 5.1 can be established using the same arguments as in the proofs of Theorem 1.2 and Proposition 3.6. ∎